1/6(6a+18)
3a-6(a-1)
3(4y-2)-(3y-3)
0.13x + 0.08 (6000-x)

Answer by JulietG(1812)

In the future, please post problems one at a time. Thank you.


1/6(6a+18)
Distribute the 1/6 to 6a and to 18
1/6(6a+18) = 1/6*6a + 1/6*18
=a+3

3a-6(a-1)
=3a-(6*a)-(6*-1)
=3a-6a- -6
=-3a+6

3(4y-2)-(3y-3)
=(3*4y)+(3*-2)-(3y-3)
=12y+ -6 - 3y --3
=9y -6 +3
=9y-3

0.13x + 0.08 (6000-x)
=.13x + (.08*6000)-.08*x
=.13x + 480 - .08x
= .05x + 480